OpenStudy (anonymous):

Find the sum 2n^3 + 4n^2 - 7 and -n^3 - 8n - 2 A. -n^3 + 4n^2 - 8n - 9 B. n^3 + 4n^2 - 8n - 9 C. n^3 + 4n^2 - 8n - 5 D. n^3 - 4n^2 -9

OpenStudy (michele_laino):

hint: \[\begin{gathered} \left( {2{n^3} + 4{n^2} - 7} \right) + \left( { - {n^3} - 8n - 2} \right) = \hfill \\ = 2{n^3} - {n^3} + 4{n^2} - 8n - 7 - 2 = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

n^3 + 4n^2 - 8n - 9

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

Thanks for the help! =)

OpenStudy (michele_laino):

Thanks! :)